Q2/// Find Thevenins equivalent across the 2 Ohm resistor for the *.circuit below 20 T lov 2V 42 22 -6v 8V 332 34

The question of electricity needs half an hour
Answer #1

not mentioned, I am assuming right to left. To find Virthremove 22 resistor, the direction is + 4V 2V + + mw or 402 + + lov V-2+V7-8-312-3(1, -3,)-620 -16+ VT -3 (2) -3(-3) = 0 Vm = 16+ 6-9 = 13 V :: Von = 13v (with polarity - +) RIL OY Yn = -13v (wiRm = 2.4642852 Rm IL M 2-4642852 RL mm 22 + V. = 13V I= Um/RmtRL - 2-46428572 PL = 2:912 A (a) - Vth = 13v, Rom - 2.4642851